# ManiSkill3: GPU Parallelized Robotics Simulation and Rendering for Generalizable Embodied AI

### v1

@@ Abstract
Simulation has enabled unprecedented compute-scalable approaches to robot learning. However, many existing simulation frameworks typically support a narrow range of scenes/tasks and lack features critical for scaling generalizable robotics and sim2real. We introduce and open source ManiSkill3, the fastest state-visual GPU parallelized robotics simulator with contact-rich physics targeting generalizable manipulation. ManiSkill3 supports GPU parallelization of many aspects including simulation+rendering, heterogeneous simulation, pointclouds/voxels visual input, and more. Simulation with rendering on ManiSkill3 can run 10-1000x faster with 2-3x less GPU memory usage than other platforms, achieving up to 30,000+ FPS in benchmarked environments due to minimal python/pytorch overhead in the system, simulation on the GPU, and the use of the SAPIEN parallel rendering system. Tasks that used to take hours to train can now take minutes. We further provide the most comprehensive range of GPU parallelized environments/tasks spanning 12 distinct domains including but not limited to mobile manipulation for tasks such as drawing, humanoids, and dextrous manipulation in realistic scenes designed by artists or real-world digital twins. In addition, millions of demonstration frames are provided from motion planning, RL, and teleoperation. ManiSkill3 also provides a comprehensive set of baselines that span popular RL and learning-from-demonstrations algorithms. Website: http://maniskill.ai/
@@
Figure 1: Multiple distinct task categories are displayed, ranging from room-scale tasks to humanoid interactions and drawing tasks. Every task shown is GPU-parallelized, simulating + rendering at state-of-the-art speeds and GPU memory efficiency. Scenes are from ReplicaCAD and AI2-THOR.
1 Introduction
@@
One of the grand challenges of robotics is robust and generalized manipulation. However, unlike vision and language research, there are still no good datasets for robotic manipulation that can be trained on. One approach has been to create human-scalable real-world teleoperation tools (Fu et al., 2024; Cheng et al., 2024) to then perform imitation learning. Another is to set up real-world reinforcement learning to fine-tune trained offline policies (Feng et al., 2023). However, real-world imitation learning approaches require enormous amounts of data that are infeasible to collect efficiently at low costs only to achieve relatively low success rates that are otherwise impractical for real-world deployment (Zhao et al., 2024). Real-world reinforcement learning approaches are promising, but require extensive setup in the real world to generate real-world rewards/success and environment resets.
@@
GPU parallelized simulations such as Isaac (Makoviychuk et al., 2021) and Mujoco’s MJX (Todorov et al., 2012) have made massive advancements in solving some robotics problems such as robot locomotion by training in large-scale GPU parallelized simulations with reinforcement learning (RL) (Rudin et al., 2021). GPU parallelized simulation makes data incredibly cheap to generate. However, when it comes to manipulation, success is often limited to narrower ranges of manipulation tasks and typically requires strong state estimation (Handa et al., 2023) to replace visual inputs like RGB or pointcloud. Existing GPU simulators have limitations that hinder the generalization and scalability of previous work. These simulators lack support for heterogeneous simulation, where each parallel environment contains different scenes. Additionally, they often don’t support fast parallel rendering capabilities. As a result, algorithms like reinforcement learning (RL) that operate on visual input train too slowly to be practical.
@@
1) State-of-the-art GPU Parallelized Simulation and Rendering: RL algorithms like PPO (Schulman et al., 2017) can now solve visual tasks in 10-1000x less time it would have taken on other simulators due to fast parallel rendering and low overhead in the system design of ManiSkill3, leading to highly efficient use of the GPU. Depending on task the simulation + rendering FPS can reach up to 30,000+, massively accelerating visual data collection by 10-1000x compared to other simulators. Importantly ManiSkill3 maintains extremely low GPU memory usage, typically 2-3x lower than that of other simulators which enables on device visual RL and larger neural networks during training.
2) Most comprehensive range of environments with 12 different categories of environments and 20+ different robots provided out of the box, all GPU parallelized: ManiSkill3 out of the box provides a diverse set of different types of environments including but not limited to mobile manipulation, room-scale scenes, drawing, and humanoid/bi-manual manipulation. We further support 20+ different robot embodiments out of the box such as quadrupeds, floating grippers, humanoids, and dextrous hands. Furthermore we support several sim2real and real2sim setups for manipulation. Importantly, extensive documentation/tutorials are provided to teach users on how to add new environments/robots, as well as how to make open-source contributions to expand the repository of simulated tasks/robots.
@@
4) Simple Unified API to Easily Manage and Build GPU Simulated Tasks: ManiSkill3 distinguishes itself from other GPU-parallelized robotics simulators and benchmarks by offering a user-friendly API for creating diverse robotics environments. Key improvements include object-oriented APIs and the elimination of complex tensor indexing. The platform provides feature-rich tooling to streamline various operations, such as domain randomization (e.g., camera poses, robot controllers), trajectory replay, action space conversion, and more.
@@ 2 Related Work
GPU Parallelized Simulation
GPU Parallelized Rendering
Heterogeneous GPU Parallelized Simulation
@@
Table 1: Comparison of major features across different open-source robotics frameworks/tools
@@
Robotics Simulation Frameworks: Isaac Lab (previously called Isaac Orbit) (Mittal et al., 2023) and Mujoco (Todorov et al., 2012) are some open-source general-purpose rigid-body GPU parallelized robotics simulators. Isaac Lab and Brax (Freeman et al., 2021) (which supports the Mujoco MJX backend) are the most similar to ManiSkill in that they provides out of the box environments for reinforcement learning/imitation learning, as well as APIs to build environments. There are robotics frameworks like Robocasa (Nasiriany et al., 2024), Habitat, (Szot et al., 2021), AI2THOR (Kolve et al., 2017), OmniGibson (Li et al., 2022), RLBench (James et al., 2020) that only have CPU simulation backends and thus run magnitudes slower than Isaac Lab and ManiSkill, limiting researchers to often only explore imitation learning/motion planning approaches instead of reinforcement learning/online learning from demonstrations methods. Isaac Lab relies on the closed source Isaac Sim framework for GPU parallelized simulation and rendering whereas ManiSkill relies on the open-source SAPIEN (Xiang et al., 2020) for the same features. Brax/Mujoco uses the MJX backend and currently does not have parallel rendering. Both Isaac Lab and ManiSkill use PhysX for GPU simulation.
@@
Robotics Datasets: Amongst existing datasets there are typically two kinds, real-world and simulated datasets. Open-X (Collaboration et al., 2023) is one of the largest real-world robotics datasets but suffers from issues with inconsistent data labels and overall poor data quality. DROID (Khazatsky et al., 2024) addresses some of Open-X’s problems by using a consistant data collection platform. However, both Open-X and DROID require immense amounts of human labor to collect data and are inherently difficult to scale up to the sizes of typical vision/language datasets. Among simulated datasets, frameworks like AI2-THOR (Kolve et al., 2017), and OmniGibson (Li et al., 2022) have complex room-scale scenes but do not readily provide demonstrations or ways to generate large-scale demonstrations for use in robot learning. Robocasa has a myriad of tasks and realistic room-scale scenes, but further leverages MimicGen (Mandlekar et al., 2023) to scale human teleoperated demonstrations by generating new demonstrations.
@@
ManiSkill sources large-scale demonstrations through a combination of different methods. For easier tasks, motion planning and rewards for RL are used to generate demonstrations. For more complex tasks without easily defined motion planning scripts or reward functions, ManiSkill3 relies on online learning from demonstrations algorithms like RLPD (Ball et al., 2023) and RFCL (Tao et al., 2024), which are more flexible compared to MimicGen used in Robocasa as MimicGen makes a number of assumptions about the task (end-effector action spaces, little to no geometric variations, engineered task stage indicators).
3 Core Features of ManiSkill3
@@
ManiSkill3 is the most feature-rich GPU simulation framework compared to popular alternatives as shown in Table 1. For the largest features, we detail them in subsequent subsections below.
3.1 Unified GPU Parallelized Tasks Supported Out Of The Box
@@
The engineering of ManiSkill3 enables it to easily support many different kinds of task categories via a flexible task-building API, contributing to a unified interface for GPU parallelized robotics simulation. Of the existing popular robotics simulators ManiSkill3 supports the most categories of different tasks. Concretely we categorize the 12 distinct categories as follows: Table top manipulation, mobile manipulation, room-scale scenes for manipulation, quadruped/humanoid locomotion, humanoid/bi-manual manipulation, multi-agent robotics, drawing/cleaning, dextrous manipulation, vision-tactile manipulation, classic control, digital twins, and soft body manipulation environments. All of these tasks are GPU parallelized and can be rendered fast in parallel as well, with examples of the tasks shown in Fig. 1. Each of these task categories has various optimizations done to run more accurately and faster to achieve state-of-the-art simulation speeds. Other simulators typically support a smaller subset of the type of tasks ManiSkill3 supports easily. Additional details on the exact optimizations/implementations and available robots are detailed in Appendix A.
3.2 GPU Parallelized Simulation and Rendering
Figure 2: GPU Simulation+Rendering speeds of various tasks with a single 128x128 resolution camera that renders every 2 physics steps in the simulation. RGB, depth, and segmentation data are all simultaneously being rendered. The only big variations the between environments of the three curves are the objects and robots being simulated.
ManiSkill3 distinguishes itself from its predecessors and other robotics simulators by offering robust support for GPU-parallelized simulation and rendering. ManiSkill3 is the first benchmark to enable fast RL from visual inputs on complex robot manipulation tasks, with Isaac Lab recently adding a similar feature. Tasks such as picking up a cube or controlling a quadruped to reach a goal from pixel inputs are now solved on the order of minutes instead of hours. RL training results/speed are detailed in Section 4.3.
@@ 3.2 GPU Parallelized Simulation and Rendering
Figure 3: GPU Simulation+Rendering of RGB/Depth speeds of the Cartpole environment with a 640x480 resolution camera in ManiSkill3 and Isaac Lab. Annotated numbers indicate GPU memory usage, with no data points beyond 128 environments for Isaac Lab due to running out of GPU memory. Note that this rendering setting mimics that of real world datasets collected in Open-X. Speed is dependent on a few factors, primarily the number of objects, geometry complexity of each object, as well as simulation/rendering configurations which can be tuned for speed or accuracy. As a result, it is possible the numbers/trends here may not hold for every environment.
@@
The performance results shown in Figure 2 are the results after simulating + rendering RGB, depth, and segmentation data simultaneously for various tasks. In terms of speed and GPU memory use, Figure 3 shows ManiSkill3 outperforms Isaac Lab, particularly when it comes to rendering common real-world camera resolutions which are important for sim2real and real2sim transfer. We compare against the latest Isaac Lab v1.2.0 released on Sept. 20, 2024. In particular, with 128 parallel environments for the benchmarked task, ManiSkill3 uses just 4.4GB of GPU memory whereas Isaac Lab uses 14.1GB. The memory efficiency of the ManiSkill3 platform allows for more room for e.g. RL replay buffers or larger neural network models. Training and inference can be kept extremely optimized on a single GPU as a result without needing to store any data on the CPU. From experimentation with visual RL, we find that GPU memory efficiency becomes much more important as the FPS gains from more parallel environments become marginal. GPU memory efficiency is especially important for off-policy algorithms like TD-MPC2 (Hansen et al., 2024) and SAC (Haarnoja et al., 2018) that typically maintain replay buffer sizes on the order of $10^{5}\sim 10^{6}$ frames. For more in-depth performance benchmarking results and comparisons of rendered outputs, see Appendix B. We acknowledge that this comparison is not strictly apples-to-apples due to differences in rendering techniques. Isaac Lab employs ray-tracing for parallel rendering, while the ManiSkill3 results are generated using SAPIEN’s rasterization renderer (see Figure 4 for a visual comparison), although ManiSkill3 also supports a ray-tracing mode without parallelization. Ray-tracing generally offers greater flexibility in balancing rendering speed and quality through the adjustment of parameters such as samples per pixel. It’s worth noting that the Isaac Lab data presented here uses the fastest rendering settings, although it can be easily tuned to achieve better rendering quality that may be helpful for sim2real. Despite the use of different rendering techniques, we believe this experiment provides a meaningful basis for comparison.
@@
GPU parallelized simulation and rendering further enable an entirely new regime of possible domain randomizations. For example you can now quickly render 1000s of different cameras, each with different extrinsics/intrinsics, mounted/fixed, as well as randomize object textures in each of the parallel environments. A subset of 4 out of 1024 environments is shown rendered by our parallel renderer in Figure 5 with different settings. This type of visual diversity in simulation data generation enables much faster training of more visually robust policies and is critical for sim2real applications. Furthermore, ManiSkill3 supports parallelized rendering of voxel and pointcloud formats to support 3D robot learning approaches.
Finally, ManiSkill3 enables extremely fast visual digital twins. For example, ManiSkill3 implements 4 of the environments in SIMPLER (Li et al., 2024) which are evaluation digital twins that enable the evaluation of generalist robotic policies trained on real-world data like Octo (Octo Model Team et al., 2024) and Robotics Transformers (RT) (Brohan et al., 2022). Currently, ManiSkill3 digital twins can evaluate models like Octo at 60x to 100x the speed of the real world without human supervision and runs approximately 10x faster than the original digital twin implementations in SIMPLER. The speed increase is due to fast and efficient parallel rendering of large camera resolutions (640x480) and flexible GPU parallelized controllers to match most real-world robot arms/manipulators. More details on sim2real and real2sim support are covered in Section 3.4.
3.3 Heterogeneous GPU Simulation

### v2

# Demonstrating GPU Parallelized Robot Simulation and Rendering for Generalizable Embodied AI with ManiSkill3

@@ Abstract
Simulation has enabled unprecedented compute-scalable approaches to robot learning. However, many existing simulation frameworks typically support a narrow range of scenes/tasks and lack features critical for scaling generalizable robotics and sim2real. We introduce and open source ManiSkill3, the fastest state-visual GPU parallelized robotics simulator with contact-rich physics targeting generalizable manipulation. ManiSkill3 supports GPU parallelization of many aspects including simulation+rendering, heterogeneous simulation, pointclouds/voxels visual input, and more. GPU Simulation with rendering on ManiSkill3 uses 2-3x less GPU memory usage than other platforms and achieves up to 30,000+ FPS in benchmarked environments due to minimal python/pytorch overhead in the system, simulation on the GPU, and the use of the SAPIEN parallel rendering system. Tasks that used to take hours to train can now take minutes. We further provide the most comprehensive range of GPU parallelized environments/tasks spanning 12 distinct domains including but not limited to mobile manipulation, drawing, humanoids, and dextrous manipulation in realistic scenes designed by artists or real-world digital twins. In addition, millions of demonstration frames are provided from motion planning, RL, and teleoperation. ManiSkill3 also provides a comprehensive set of baselines that span popular RL and learning-from-demonstrations algorithms.
@@
Figure 1: Multiple distinct task categories are displayed, ranging from room-scale tasks to humanoid interactions and drawing tasks. Majority of tasks shown are GPU-parallelized, simulating + rendering at state-of-the-art speeds and GPU memory efficiency. Scenes are from ReplicaCAD and AI2-THOR.
I Introduction
@@
One of the grand challenges of robotics is robust and generalized manipulation. However, unlike vision and language research, there are still no good datasets for robotic manipulation that can be trained on. One approach has been to create human-scalable real-world teleoperation tools [18, 12] to then perform imitation learning. Another is to set up real-world reinforcement learning to fine-tune offline trained policies [16]. However, real-world imitation learning approaches require enormous amounts of data that are infeasible to collect efficiently at low costs only to achieve relatively low success rates that are otherwise impractical for real-world deployment [54]. Real-world reinforcement learning approaches are promising, but require extensive setup in the real world to generate real-world rewards/success and environment resets.
@@
GPU parallelized simulations such as Isaac [33] and Mujoco’s MJX [49] have made massive advancements in solving some robotics problems such as robot locomotion by training in large-scale GPU parallelized simulations with reinforcement learning (RL) [43]. GPU parallelized simulation makes data incredibly cheap to generate. However, when it comes to manipulation, success is often limited to narrower ranges of manipulation tasks and typically requires strong state estimation [21] to replace visual inputs like RGB or pointcloud. Existing GPU simulators have limitations that hinder the generalization and scalability of previous work. These simulators lack support for heterogeneous simulation, where each parallel environment contains different scenes. Additionally, they often don’t support fast parallel rendering capabilities. As a result, algorithms like reinforcement learning (RL) that operate on visual input train too slowly to be practical. We propose ManiSkill3 to address past limitations and open source the framework under the Apache-2.0 license, building upon past work in ManiSkill 1 and 2 [38, 19]. The majority of features presented in this paper are already open-sourced. Website: maniskill.ai/. Documentation/Tutorials: maniskill.readthedocs.io/. Video demo gallery: maniskill.readthedocs.io/en/latest/user_guide/demos/gallery.html.
@@
1) State-of-the-art GPU Parallelized Simulation and Rendering: RL algorithms like PPO [44] can now solve visual tasks faster than it would have taken on other simulators due to fast parallel rendering and low overhead in the system design of ManiSkill3, leading to highly efficient use of the GPU. Depending on task the simulation + rendering FPS can reach up to 30,000+, massively accelerating visual data collection relative to most other simulators. Importantly ManiSkill3 maintains extremely low GPU memory usage, typically 2-3x lower than that of other simulators which enables on device visual RL and larger neural networks during training.
2) Most comprehensive range of environments with 12 different categories of environments and 20+ different robots provided out of the box: ManiSkill3 out of the box provides a diverse set of different types of environments including but not limited to mobile manipulation, room-scale scenes, drawing, and humanoid/bi-manual manipulation. We further support 20+ different robot embodiments out of the box such as quadrupeds, floating grippers, humanoids, and dextrous hands. Furthermore we support several sim2real and real2sim setups for manipulation. Importantly, extensive documentation/tutorials are provided to teach users on how to add new environments/robots, as well as how to make open-source contributions to expand the repository of simulated tasks/robots. A core focus of ManiSkill3 is not on building a lot of environments in each category, but instead building many template/examples that users can then build on top of themselves for their own use-cases.
@@
4) Simple Unified API to Easily Manage and Build GPU Simulated Tasks: ManiSkill3 distinguishes itself from other GPU-parallelized robotics simulators and benchmarks by offering a user-friendly API for creating diverse robotics environments. Improvements include object-oriented APIs and the elimination of complex tensor indexing. The platform provides feature-rich tooling to streamline various operations, such as domain randomization (e.g., camera poses, object materials), trajectory replay, controller action conversion, and more.
@@ II Related Work
Parallelized Simulation
Parallelized Rendering
Parallelized Heterogeneous Simulation
@@
TABLE I: Comparison of major features across different open-source robotics frameworks/tools.
@@
Robotics Simulation Frameworks: Isaac Lab (previously called Isaac Orbit) [36] and Mujoco [49] are some open-source general-purpose rigid-body GPU parallelized robotics simulators. Isaac Lab and Brax [17] (which supports the Mujoco MJX backend) are the most similar to ManiSkill3 in that they provides out of the box environments for reinforcement learning/imitation learning, as well as APIs to build environments. There are robotics frameworks like Robocasa [39], Habitat, [46], AI2THOR [29], OmniGibson [30], RLBench [25] that only have CPU sim backends and thus run magnitudes slower than Isaac Lab/ManiSkill3, limiting researchers to often only explore imitation learning/motion planning approaches instead of reinforcement learning/online learning from demonstrations methods. Isaac Lab relies on the closed source Isaac Sim framework for GPU parallelized simulation and rendering whereas ManiSkill3 relies on the open-source SAPIEN [52] for the same features. Brax/Mujoco uses the MJX backend and currently does not have parallel rendering. Both Isaac Lab and ManiSkill3 use PhysX for GPU simulation. Note that the environments and baselines leveraging the fast parallel rendering in ManiSkill3 is concurrent work to Isaac Lab.
@@
Robotics Datasets: Amongst existing datasets there are typically two kinds, real-world and simulated datasets. Open-X [14] is one of the largest real-world robotics datasets but suffers from issues with inconsistent data labels and overall poor data quality. DROID [28] addresses some of Open-X’s problems by using a consistant data collection platform. However, both Open-X and DROID require immense amounts of human labor to collect data and are inherently difficult to scale up to the sizes of typical vision/language datasets. Among simulated datasets, frameworks like AI2-THOR [29], and OmniGibson [30] have complex room-scale scenes but do not readily provide demonstrations or ways to generate large-scale demonstrations for use in robot learning. Robocasa has a myriad of tasks and realistic room-scale scenes, but further leverages MimicGen [35] to scale human teleoperated demonstrations by generating new demonstrations.
@@
ManiSkill3 sources large-scale demonstrations through a combination of different methods. For easier tasks, motion planning and rewards for RL are used to generate demonstrations. For more complex tasks without easily defined motion planning scripts or reward functions, ManiSkill3 relies on online learning from demonstrations algorithms like RLPD [2] and RFCL [47], which are more flexible compared to MimicGen used in Robocasa as MimicGen makes a number of assumptions about the task (end-effector action spaces, little to no geometric variations, engineered task stage indicators).
III Core Features of ManiSkill3
@@
ManiSkill3 is the most feature-rich GPU simulation framework compared to popular alternatives as shown in Table I. For the largest features, we detail them in the subsections below.
III-A Diverse Tasks Supported Out Of The Box
@@
The design of ManiSkill3 enables support for many different kinds of task categories via a flexible task-building API. Of the existing popular robotics simulators ManiSkill3 supports the most categories of different tasks. Concretely we categorize the 12 distinct categories as follows: Table top manipulation, mobile manipulation, room-scale scenes for manipulation, quadruped/humanoid locomotion, humanoid/bi-manual manipulation, multi-agent robotics, drawing/cleaning, dextrous manipulation, vision-tactile manipulation, classic control, digital twins, and soft body manipulation environments. The majority of these tasks are GPU parallelized and can be rendered fast in parallel as well, with examples of the tasks shown in Fig. 1. Each of these task categories have various optimizations done to run more accurately and/or faster. Other simulators typically support a smaller subset of the type of tasks ManiSkill3 supports easily. Additional details on the exact optimizations/implementations and available robots are detailed in Appendix VII.
@@
Figure 2: GPU Simulation+Rendering of RGB speeds of the Cartpole environment with different camera setups ManiSkill3 and Isaac Lab. Annotated numbers indicate GPU memory usage, with no data points beyond 128 environments for Isaac Lab due to running out of GPU memory. Note that this rendering setting mimics that of real world datasets collected in Open-X and Droid. Speed is dependent on a few factors, primarily the number of objects, geometry complexity of each object, as well as simulation/rendering configurations which can be tuned for speed or accuracy. As a result, it is possible the numbers/trends here may not hold for every environment.
@@
III-B GPU Parallelized Simulation and Rendering
Figure 4: GPU Simulation+Rendering speeds of various tasks with a single 128x128 resolution camera with a simulation frequency of 120 and control frequency of 60, meaning the camera renders ever 2 sim steps. RGB, depth, and segmentation data are all simultaneously being rendered. The only big variations between environments of the three curves are the objects and robots being simulated.
ManiSkill3 distinguishes itself from its predecessors and other robotics simulators by offering robust support for GPU-parallelized simulation and rendering. ManiSkill3 is the first general benchmark to enable fast RL from visual inputs on complex robot manipulation tasks, with Isaac Lab recently adding a similar feature. Tasks such as picking up a cube or controlling a quadruped to reach a goal from pixel inputs are now solved on the order of minutes instead of hours. RL training results/speed are detailed in Section IV-C.
The performance results shown in Figure 4 are the results after simulating + rendering RGB, depth, and segmentation data simultaneously for various tasks. In terms of speed and GPU memory use, Figure 2 shows ManiSkill3 outperforms Isaac Lab, particularly when it comes to rendering common real-world camera resolutions which can be important for sim2real and real2sim transfer. In particular, with 128 parallel environments for the benchmarked task, ManiSkill3 uses just 3.5GB of GPU memory whereas Isaac Lab uses 14.1GB. The memory efficiency of the ManiSkill3 platform allows for more room for e.g. RL replay buffers or larger neural network models such as large vision language action models. Training and inference can be kept extremely optimized on a single GPU as a result without needing to store any data on the CPU. From experimentation with visual RL, we find that GPU memory efficiency becomes much more important as the FPS gains from more parallel environments become marginal. GPU memory efficiency is especially important for off-policy algorithms like TD-MPC2 [22] and SAC [20] that typically maintain replay buffer sizes on the order of $10^{5}\sim 10^{6}$ frames. For example storing RGB data from one 128x128 camera would require at least 9GB of GPU memory for a replay buffer of size $200,000$, which can easily lead to out of memory issues. For more in-depth performance benchmarking results and comparisons of rendered outputs, see Appendix XI.
We acknowledge that this comparison is not strictly apples-to-apples due to differences in rendering techniques. Isaac Lab employs ray-tracing for parallel rendering, while the ManiSkill3 results are generated using SAPIEN’s rasterization renderer (see Figure 3 for a visual comparison), although ManiSkill3 also supports a ray-tracing mode without parallelization. Ray-tracing generally offers greater flexibility in balancing rendering speed and quality through the adjustment of parameters such as samples per pixel. It’s worth noting that the Isaac Lab data presented here uses the fastest rendering settings, although it can be easily tuned to achieve better rendering quality that may be helpful for sim2real. Despite the use of different rendering techniques, we believe this experiment provides a meaningful basis for comparison.
@@ III-B GPU Parallelized Simulation and Rendering
GPU parallelized simulation and rendering enable an entirely new regime of running efficient domain randomizations. For example you can now quickly render over a 1000 different cameras, each with different extrinsics/intrinsics, mounted/fixed, as well as randomize object textures in each of the parallel environments. A subset of 4 out of 1024 environments renders are shown in Figure 5 with different settings. This type of visual diversity in simulation enables much faster training of more visually robust policies and is critical for sim2real applications. Furthermore, ManiSkill3 supports parallelized rendering of voxel and pointcloud formats necessary for 3D robot learning approaches [53, 24, 45].
Finally, ManiSkill3 enables extremely fast visual digital twins. For example, ManiSkill3 implements 4 of the environments in SIMPLER [31] which are evaluation digital twins that enable the evaluation of generalist robotic policies trained on real-world data like Octo [40] and Robotics Transformers (RT) [3]. ManiSkill3 digital twins can evaluate models like Octo at 60x to 100x the speed of the real world without human supervision, approximately 10x faster than the original digital twin implementations in SIMPLER. The speed increase is due to fast and efficient parallel rendering of large camera resolutions (640x480) and flexible GPU parallelized controllers to match most real-world robot arms/manipulators. More details on sim2real/real2sim support are covered in Section III-E.
III-C Heterogeneous GPU Simulation